OpenStudy (anonymous):

Solve the system: x+y+z=4 x+2y-z=2 3z=x-3y

OpenStudy (luigi0210):

matrix time

OpenStudy (loser66):

100% agree

OpenStudy (luigi0210):

set it up as a matrix

OpenStudy (texaschic101):

or elimination

OpenStudy (loser66):

@texaschic101 not easy to go on that way

OpenStudy (anonymous):

how would you use elimination for this problem?

OpenStudy (loser66):

you are better set up matrix, texaschic off line. cannot guide you

OpenStudy (whpalmer4):

Not really that hard with elimination. \[x+y+z=4\]\[x+2y-z=2\]\[-x+3y +3z= 0\] Add rows 1 and 2 to get \[2x+3y = 6\] Add rows 2 (mult by 3) and 3 to get \[2x+9y = 6\]Subtract the two resulting rows to get an equation that is very easy to solve :-) Plug result into previous equations.

OpenStudy (texaschic101):

I prefer elimination....that is just my opinion